APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday August 30, 2007 (MJD 54343) ---=== OBSERVING TEAM ===--- dayShift: Howard Brewington Observers: Elena Malanushenko, Steph Snedden ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- More high humidity. Never opened. A good night for mildew. ---=== OBSERVING LOG ===--- Afternoon: ---------- The spectro focus/collimation was checked but not adjusted. The focus is set for about 14-degrees C. Watcher reported sp2's mechanicals as, "Not Connected." Other misc Watcher messages are posted in the Problems Section. 00:00Z Opened louvers, cracked doors, turned on fans. 01:00Z Closed louvers and doors and turned fans off because of high humidity. Night: ---------- High humidity persisted. There were a few bounces to lower humidity, and we almost opened twice, but the humidity went back up. Water started dripping off the eaves of the buildings. That clinched it for the night. 11:00Z - We closed down. endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6911 22:06Z 22:50Z 100 O 19.52 30.49 87 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 46181, plate -9999 ------- 22:07 46181 0.0 bias ----- sequence 46182, plate 2459 ------- 22:23 46182 10.0 flat spectro focus 22:26 46183 2.0 arc " " 22:30 46184 3.0 arc " " 22:32 46185 3.0 arc " " ----- sequence 46186, plate -9999 ------- 01:12 46186 0.0 bias ----- sequence 46187, plate 2459 ------- 01:29 46187 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:03Z 59 54 0 321 (NW) 0 - partly cloudy 21:35Z 59 53 2 227 (SW) 0 - ' 22:07Z 57 53 3 277 (W) 0 - ' 22:37Z 56 52 5 1 (N) 0 - ' 23:08Z 58 53 0 153 (SSE) 0 - ' 23:38Z 59 48 2 164 (SSE) 0 - ' 00:08Z 58 48 7 100 (E) 0 - ' 00:38Z 57 48 9 83 (E) 0 - ' 01:19Z 55 52 9 87 (E) 0 - ' 01:53Z 54 51 9 85 (E) 0 - mostly clear 02:25Z 53 52 12 76 (ENE) 0 - ' 02:57Z 53 51 7 91 (E) 0 - ' 03:29Z 53 50 6 88 (E) 0 - ' 04:01Z 53 49 6 86 (E) 0 - ' 04:33Z 53 49 8 81 (E) 0 - ' 05:05Z 53 48 9 87 (E) 0 - partly cloudy 05:37Z 51 50 11 77 (ENE) 0 - ' 06:10Z 53 48 12 77 (ENE) 0 - ' 06:42Z 52 49 12 69 (ENE) 0 - variably cloudy 07:14Z 50 49 13 76 (ENE) 0 - ' 07:46Z 50 49 12 40 (NE) 0 - ' 08:18Z 54 46 10 33 (NNE) 0 - ' 08:51Z 53 48 8 14 (NNE) 0 - ' 09:23Z 54 48 11 30 (NNE) 0 - ' 09:55Z 54 48 10 30 (NNE) 0 - ' 10:27Z 54 47 6 3 (N) 0 - ' 10:59Z 53 47 6 31 (NNE) 0 - ' ---=== TELESCOPE STATUS ===--- 00:00Z Opened louvers, cracked doors, turned on fans. 01:00Z Ventilation back off. Doors closed 06:25Z Ventilation started, doors opened 06:45Z Ventilation off, doors closed 08:27Z Ventilation on, doors open 10:24Z Ventilation off, doors closed. 10:46Z Telescope stowed at 30 deg alt with cartridge 4 mounted. Counterweights at 275. Spectro autofill connected. 10:48Z - Spectro dewar weight: 205 lb; 21 psi 10:48Z - Imager dewar weight: 162 lb; 17 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.9.1 2007-06-21 sdssProcedures: v2_35 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2120 -5.2130 Axial B -0.2135 -0.2135 Axial C -0.0014 -0.0016 Trans D 11.6370 11.6400 Lateral E 9.9568 9.9568 Lateral F 11.5316 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 14.1 Alt.: 30.000681 MIGS TONIGHT NOMINAL Axial A 0.7974 0.7973 Axial B 0.6040 0.6041 Axial C 0.5425 0.5425 Trans D 0.5996 0.5999 GALILS Commanded: 1799888. 1309521. 1643321. -9450. -4400. Actual: 1798650. 1309550. 1643400. -9450. -4400. SETMIR VALUES SecDesOrient: 1257.00 -60.00 -140.00 -190.00 200.12 SecOrient: 1257.31 -59.83 -140.43 -189.75 200.02 ---=== PROBLEMS IN DETAIL ===--- --== Misc Watcher Messages ==-- The following errors appear to be related to a communications timeout. Note they all have the same timestamp. m1 temp (Excessive M1 - PMSS Difference:Excessive Telescope - Ambient Difference:Excessive Radial Gradient:Excessive Front-Back Difference (Entire Mirror):Excessive Front-Back Difference (Left Half):Excessive Front-Back Difference (Right Half):Excessive Front-Back Difference (Lower Half):Excessive Front-Back Difference (Upper Half): 2007-08-30 22:10:50Z ) m1 mig (Communications Fault: 2007-08-30 22:10:50Z) tmicro (Disconnected Sensor in M2B7:Disconnected Sensor in M2B6:Disconnected Sensor in M2B5:Disconnected Sensor in M2B4:Disconnected Sensor in M2B3:Disconnected Sensor in M2B2:Disconnected Sensor in M2B1:Disconnected Sensor in M2B0:Disconnected Sensor in M1B7:Disconnected Sensor in M1B6:Disconnected Sensor in M1B5:Disconnected Sensor in M1B4:Disconnected Sensor in M1B3:Disconnected Sensor in M1B2:Disconnected Sensor in M1B1:Disconnected Sensor in M1B0:Disconnected Sensor in M0B7:Disconnected Sensor in M0B6:Disconnected Sensor in M0B5:Disconnected Sensor in M0B4:Disconnected Sensor in M0B3:Disconnected Sensor in M0B2:Disconnected Sensor in M0B1:Disconnected Sensor in M0B0:Communications Fault: 2007-08-30 22:10:50Z ) m2 galil (Communications Fault: 2007-08-30 22:10:50Z) m1 galil (Communications Fault: 2007-08-30 22:10:50Z) m2 mig (Communications Fault: 2007-08-30 22:10:50Z) --== Spectro Focus Numbers ==-- IDL 5.5a> collimate, 46184 Reading /data/spectro/54343/sdR-b1-00046184.fit Reading /data/spectro/54343/sdR-b1-00046185.fit COLLIMATE: Min offset = -0.81 pix COLLIMATE: Max offset = 0.65 pix COLLIMATE: RMS across CCD = 0.32 pix COLLIMATE: Mean offset = -0.04 pix COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 1.4 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54343-b1-00046184.log PostScript file = Collimate-54343-b1-00046184.ps Reading /data/spectro/54343/sdR-b2-00046184.fit Reading /data/spectro/54343/sdR-b2-00046185.fit COLLIMATE: Min offset = -0.10 pix COLLIMATE: Max offset = 0.47 pix COLLIMATE: RMS across CCD = 0.15 pix COLLIMATE: Mean offset = 0.24 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -8.4 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54343-b2-00046184.log PostScript file = Collimate-54343-b2-00046184.ps Reading /data/spectro/54343/sdR-r1-00046184.fit Reading /data/spectro/54343/sdR-r1-00046185.fit COLLIMATE: Min offset = -0.58 pix COLLIMATE: Max offset = 0.61 pix COLLIMATE: RMS across CCD = 0.29 pix COLLIMATE: Mean offset = -0.07 pix COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 651 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 651" Log file = Collimate-54343-r1-00046184.log PostScript file = Collimate-54343-r1-00046184.ps Reading /data/spectro/54343/sdR-r2-00046184.fit Reading /data/spectro/54343/sdR-r2-00046185.fit COLLIMATE: Min offset = -0.80 pix COLLIMATE: Max offset = 0.18 pix COLLIMATE: RMS across CCD = 0.25 pix COLLIMATE: Mean offset = -0.16 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 1478 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 1478" Log file = Collimate-54343-r2-00046184.log PostScript file = Collimate-54343-r2-00046184.ps % Program caused arithmetic error: Floating illegal operand IDL 5.5a>